OpenStudy (aaronandyson):

two cars left place A and reached place B simultaneously in 2hr. the first car traveled half the distance with 30km/hr and the other half at the speed of 45 km/hr .the second car at the same time covered the entire with a constant acceleration starting from rest.find the acceleration of the second car

OpenStudy (aaronandyson):

@mayankdevnani samjha bhai muhje yeh sum...

OpenStudy (mayankdevnani):

apna work dikha bhai? dimag kaam nahi kr raha h

OpenStudy (mayankdevnani):

kuch galat huya toh bata dunga

OpenStudy (aaronandyson):

avg vel nikaloon?

OpenStudy (mayankdevnani):

nahi uski jarurat nahi h

OpenStudy (aaronandyson):

total vel 75 ms h na?

OpenStudy (mayankdevnani):

haan dekh le avg vel use kr le

OpenStudy (aaronandyson):

avg vel = 2*30*45 all over 30+45 = 2700/75=36ms

OpenStudy (mayankdevnani):

yeh kaise aaya? 2700 kaise aaya?

OpenStudy (aaronandyson):

i used this eqn: \[\frac{ 2v_1v_2 }{ v_1 + v_2 }\]

OpenStudy (aaronandyson):

30*45* = 1350*2 2*1350 = 2700